OpenStudy (anonymous):

write an equation in standard form for a line with a slope of -1 passing through C(2,1)

OpenStudy (texaschic101):

put it in point slope formula : y - y1 = m(x - x1) m = -1 (2,1) x1 = 2 and y1 = 1 now lets sub our info into the formula... y - 1 = -1(x - 2) y - 1 = -x + 2 y = -x + 2 + 1 y = -x + 3 <-- slope intercept form now we need to put it in standard form Ax + by = C y = -x + 3 x + y = 3 <-- standard form

OpenStudy (anonymous):

thank you!

OpenStudy (texaschic101):

no problem :)